Suppose the U.S. government exempts the first $20,000 in household income from taxation (or in other words, the tax rate is 0% on the first $20,000 of income), but other than that provides no tax exemptions. On all remaining income, the U.S. government imposes a 10% tax rate on the first $30,000 of taxable income, a 20% tax rate on the next $70,000 of taxable income, and a 30% tax rate on all remaining taxable income.
Moreover, suppose that the State of Utah also imposes an additional income tax of 5% on ALL income (i.e., first $20,000 not exempt). However, the State of Utah provides a nonrefundable tax credit of $1,200. This tax credit phases out 2 cents for each dollar that total income exceeds $40,000.
Fill in the blanks denoted by the highlighted cells. Note that extra table cells are added to facilitate calculation, but not all spaces are questions.  Calculate the effective tax rates / effective marginal tax rate with respect to TOTAL income.
Federal Taxes State Taxes
Total Income Federal Taxable Income Federal Tax Liability Federal Effective Tax Rate (Average Tax Rate with Respect to Total Income) Federal Statutory Marginal Tax Rate State Tax Before Credit State Tax Credit State Tax Liability State Average Tax Rate State Effective Marginal Tax Rate
$30,000 1 5 9 13
$60,000 2 6 10 14
$90,000 3 7 11 15
$150,000 4 8 12 16
Answer #1
A B=A-20000 C D=(C/A)*100% E=5%*A F=2%*(A-40000) if A>40000 G=E-F H=(G/A)*100% I
Federal Taxes State Taxes
Total Income Federal Taxable Income Federal Tax Liability Federal Effective Tax Rate (Average Tax Rate with Respect to Total Income) Federal Statutory Marginal Tax Rate State Tax Before Credit State Tax Credit State Tax Liability State Average Tax Rate State Effective Marginal Tax Rate
$30,000 $10,000 $1,000 3.3% 10.0% $1,500 $0 $1,500 5.0% 5%
$60,000 $40,000 $5,000 8.3% 20.0% $3,000 $400 $2,600 4.3% 3.0%
$90,000 $70,000 $11,000 12.2% 20.0% $4,500 $1,000 $3,500 3.9% 3.0% (3500-2600)/(90000-60000)
$150,000 $130,000 $32,000 21.3% 30.0% $7,500 $2,200 $5,300 3.5% 3.0%
Federal Tax Liability for $30000 total Income $1,000 (10%*10000)
Federal Tax Liability for $60000 total Income $5,000 10%*30000+20%*(40000-30000)
Federal Tax Liability for $90000 total Income $11,000 (5000+20%*(70000-40000)
Federal Tax Liability for $150000 total Income $32,000 (11000+20%(30000+70000-70000)+30%(150000-100000)
Marginal tax rate of state taxes for income above $40000:
For Total income of $40000, Tax=5%*40000= $2,000
For Total income of $60000, Tax Liability $2,600
Marginal Tax $600
Marginal Income=60000-40000 20000
Marginal tax rate =600/20000 3.0%